# Expanding on the BRIAR Dataset: A Comprehensive Whole Body Biometric Recognition Resource at Extreme Distances and Real-World Scenarios (Collections 1-4)

### v1

@@ 1 Introduction
Figure 1: Sample images from the BRIAR dataset. The dataset’s rich spectrum of diverse participants allows for analysis of many facets of demographic diversity. Subjects in figures have consented to appearing in publications.
1.1 Contributions
@@
The BRIAR dataset is already being utilized across a diverse range of research initiatives. Those interested in access to the BRIAR data should contact the authors, who will coordinate the request to the appropriate government contact. BRIAR BGC1-4 is complemented by additional datasets collected under related efforts [18, 16, 8, 9, 10, 22], many of which are openly available.
The remainder of this paper is organized as follows: In Section 2 we provide background information addressing the challenges and considerations in biometrics at long-range and high-altitude and a summary of related work in this topic area. Section 3 provides a detailed overview of the data collection methodology and highlights its unique elements. Section 4 includes statistical breakdowns of the most recent additions to the dataset. Section 5 details the challenges and methodology used in curating such a feature-rich dataset with emphasis on the curation pipeline, quality analysis, metadata construction, annotations, evaluation protocol design, and other associated activities. Finally, Section 6 provides discussions and concluding remarks including plans for future extensions to the ongoing program.
2 Background and Related Work
@@
There exists a small body of work prior to the BRIAR program that considers the impact of long-range imaging to the performance of facial recognition models. Such works include the studies and datasets listed in Table 1. Although these works demonstrate the need for further study, they do not provide the required experimental diversity to explore performance impacts related to clothing changes, multiple environmental conditions, sensor/optical configuration, subject demographics, or other factors that can only be teased out via a large and complex dataset.
@@
Table 1: Comparison of the BRIAR dataset to similar large-scale, long-range, and/or multimodal studies.
* Nonvisible is currently held back for future research efforts.
** The dataset includes an additional 587 distractors with indoor images only.
3 Data Collection Methods
@@
3.1 Privacy, Security, and Well-being
@@ 3.3 Field Collection
3.4 BGC4 Mock City: Hogan’s Alley
@@ 4 Dataset Summary
5 Data Curation Methodology
@@
1. Data Cleaning and Compilation (Section 5.1.1)
@@
2. Video and Image Extraction (Section 5.1.2)
@@
3. XML Data Generation (Section 5.1.3)
@@
4. Quality Assurance and Finalization (Section 5.1.4)
@@
Table 2: Overview of data curation pipeline separated into relevant tasks by data type. Each row in the table denotes a sequential step in the annotation and curation process. *subject metadata includes demographics, activities, and camera specifications **field measurements include activities, weather, CN2, camera measurements
5.1 Data Curation Pipeline
The collected data was curated such that there is no more than one video or image for each unique subject, activity, clothing set, and sensor. For each curated video a corresponding XML file was generated that describes the activity details, subject demographics, camera specifications, camera measurements, and weather/atmospheric conditions. After curation, a set of quality assurance steps were taken to improve the overall quality of the final dataset. Each category of sensor warrants a slightly different curation process and so each has its own curation pipeline. Similarities and differences between these pipelines are shown in Table 2.
5.1.1 Data Cleaning and Compilation
@@
5.1.2 Video and Image Extraction
@@ 5.1.3 XML Metadata Generation
5.1.4 Quality Assurance and Finalization
@@
5.1.5 Group Activity Curation
@@ 5.2 Annotations
5.3 Evaluation Protocol Design
@@
6 Conclusions and Future Work
@@
7 Acknowledgements
This information is not included in review submission to keep authors anonymous. Contractual funding information will be included upon acceptance.

### v2

@@ I INTRODUCTION
Figure 1: Sample images from the BRIAR dataset. Subjects in figures have consented to appearing in publications.
I-A Contributions
@@
Accurate biometric identification systems have become a vital resource supporting security and safety initiatives. Beneficial applications are wide-ranging, from combating human trafficking and terrorism to supporting smart-city infrastructure and disaster response efforts. Even so, potential misuses of identification and recognition models as well as the data which shapes them are just as wide-ranging. Oversight and control of dataset access helps to prevent these potential misuses. For the BRIAR dataset, access requires approval from the US government sponsor and a data use agreement and review by the Institutional Review Board responsible for the dataset. Those interested in access to the BRIAR data should contact the authors, who will coordinate the request to the appropriate government contact. The BRIAR dataset is already being utilized across a diverse range of research initiatives.
The remainder of this paper is organized as follows: In Section II we provide background information addressing the challenges and considerations in biometrics at long-range and high-altitude and a summary of related work in this topic area. Section III provides a detailed overview of the data collection methodology and highlights its unique elements. Section IV includes statistical breakdowns of the most recent additions to the dataset. Section V details the challenges and methodology used in curating such a feature-rich dataset with emphasis on the curation pipeline, quality analysis, metadata construction, annotations, evaluation protocol design, and other associated activities. Finally, Section VI provides discussions and concluding remarks including plans for future extensions to the ongoing program.
II BACKGROUND AND RELATED WORK
@@
BRIAR BGC1-4 is complemented by additional datasets collected under related efforts [18, 16, 8, 9, 10, 22], many of which are openly available. There also exists a small body of work prior to the BRIAR program that considers the impact of long-range imaging to the performance of facial recognition models. Such works include the studies and datasets listed in Table I. Although these works demonstrate the need for further study, they do not provide the required experimental diversity to explore performance impacts related to clothing changes, multiple environmental conditions, sensor/optical configuration, subject demographics, or other factors that can only be teased out via a large and complex dataset.
@@
TABLE I: Comparison of the BRIAR dataset to similar large-scale, long-range, and/or multimodal studies.
Nonvisible is currently held back for future research efforts.
∗∗ The dataset includes an additional 587 distractors with indoor images only.
III DATA COLLECTION METHODS
@@
III-A Privacy, Security, and Well-being
@@ III-C Field Collection
III-D BGC4 Mock City: Hogan’s Alley
@@ IV DATASET SUMMARY
V DATA CURATION METHODOLOGY
@@
1. Data Cleaning and Compilation (Section V-A1)
@@
2. Video and Image Extraction (Section V-A2)
@@
3. XML Data Generation (Section V-A3)
@@
4. Quality Assurance and Finalization (Section V-A4)
@@
TABLE II: Overview of data curation pipeline separated into relevant tasks by data type. Each row in the table denotes a sequential step in the annotation and curation process. *subject metadata includes demographics, activities, and camera specifications **field measurements include activities, weather, CN2, camera measurements
V-A Data Curation Pipeline
The collected data was curated such that there is no more than one video or image for each unique subject, activity, clothing set, and sensor. For each curated video a corresponding XML file was generated that describes the activity details, subject demographics, camera specifications, camera measurements, and weather/atmospheric conditions. After curation, a set of quality assurance steps were taken to improve the overall quality of the final dataset. Each category of sensor warrants a slightly different curation process and so each has its own curation pipeline. Similarities and differences between these pipelines are shown in Table II.
V-A1 Data Cleaning and Compilation
@@
V-A2 Video and Image Extraction
@@ V-A3 XML Metadata Generation
V-A4 Quality Assurance and Finalization
@@
V-A5 Group Activity Curation
@@ V-B Annotations
V-C Evaluation Protocol Design
@@
VI CONCLUSIONS AND FUTURE WORKS
@@
VII ACKNOWLEDGMENTS
This research is based upon work supported by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via D20202007300010. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
This research used resources from the Knowledge Discovery Infrastructure at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
Notice: This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
The authors of this paper also wish to acknowledge the contributions of:
Knowledge Discovery Infrastructure (KDI) Staff: Dallas Sacca and Ryan Tipton.
Proctors and drivers: Raymond Borges-Hink, Nancy Engle, Dale Hensley, Nikki Jones, Michael Jones, Marylin Langston, Matt Love, Amanda Mottern, Gio Pascascio, Linda Paschal, Christina Peshoff, Donna Pierce, Ryan Styles, and Lauren Torkelson.
UAS Pilots: Joe Baldwin, Kase Clapp, Dakota Haldeman, Andrew Harter, Amanda Killingsworth, Matt Larson, Genevieve Martin, Aaron O’Toole, Jason Richards, and Brad Stinson.
Health and Safety Support Staff: Margaret Smith and Miranda Liner.
ETHICAL IMPACT STATEMENT
As discussed in subsection III-A, safe and ethical collection and data management practices are the first priority for the operation of the BRIAR team.
The BRIAR program is reviewed and approved by the Central Department of Energy Institutional Review Board. The study number is 00000094. The IRB package included a protocol, consent, scripts, checklists, and surveys. A modification was submitted for any change to the research, procedures, or team members.
Each participant signed a consent form before participating. They had the opportunity to ask questions of the designated member of team responsible for the consent process. The participant was notified of how their data would be shared and were offered the opportunity to allow their photos, videos, and likeness to be used publicly by checking a box on the consent form. This was optional and the participant was notified it was their decision.
The study posed minimal risk to the participants. The primary risk to subjects participating in BRIAR collections is the risk to privacy from disclosure of personally identifiable information (PII), i.e. a subject’s appearance and likeness. To mitigate this risk, the BRIAR datasets are stored on secure, access-controlled computers in accordance with federal requirements for the protection of PII. Any sharing of the data requires approval and a signed data use agreement (DUA).
Risks to the physical safety of participants in the study were addressed as comprehensively as possible. Subjects included in outdoor collections were offered insect repellent, sunscreen, water, snacks, shade, and bathroom facilities to mitigate any risk to being outdoors. Onsite first aid care was provided at no cost to participants. All BRIAR team members were required to be certified in first aid and CPR. Participants were always accompanied by a BRIAR team member to ensure they would not enter into areas not associated with the study and in which the participant could encounter more significant risks.
Unmanned aerial systems (UAS) were operated in support of this project. A mix of fixed-wing, tethered rotary-wing, and untethered rotary-wing aircraft were used. All aircraft were operated by certified FAA Part 107 licensed flight crews in accordance with applicable regulations and documented safety protocols. Those safety protocols dictate minimum and maximum altitudes, maintenance and inspection requirements, flight procedures, crew requirements, weather restrictions, and other safety requirements in a variety of operational conditions.
Participants were compensated based on completion of indoor and outdoor activities. They received a gift card at the end of their participation. This amount was determined based on the time and effort given by the participant and was approved by the IRB to be appropriate to prevent monetary coercion.
The study did include participants associated with the employer of the BRIAR team. This is considered a vulnerable population by the IRB. The research team implemented additional measures to protect against coercion, including ensuring that individuals recruited did not report to any BRIAR study team member. The collections have included some university locations; in each case, the university was notified of the study for review.
Access to the dataset is managed by the US Government sponsor of the BRIAR program, and requires IRB review. These measures help to ensure that data use is ethical and follows US laws and IRB regulations for the protection of civil liberties.